OpenStudy (anonymous):

URGENT! Evaluate f(x)=-2x-5 for x=3

OpenStudy (anonymous):

Evaluating means replacing x=3 in f(x)

OpenStudy (usukidoll):

let x = 3 f(3) = -2(3)-5 = ?

OpenStudy (decentnabeel):

f(x)=-2x-5 for x=3 pulg the value x=3 =-2(3)-5 =-6-5 =-11

OpenStudy (anonymous):

So, if \[f(\color{red}x)=-2\color{red}x-5\] then \[f(\color{red}3)=-2\cdot \color{red}3-5\]

OpenStudy (anonymous):

\[f(\color{red}x)=-2 \cdot \color{red}3 -5=-6-5=-11\]

OpenStudy (usukidoll):

I got -11 too f(3) = -6-5 = -11

OpenStudy (anonymous):

Thanks you guys!!